# Learning dynamical systems from data: Gradient-based dictionary optimization

### v1

@@ 4.1.3 Triple-well 2D
The potential is visualized in Figure 4a. We set $\beta=1.68$ and uniformly sample $m=100,000$ training data points in the domain $[-2,2]\times[-2,2]$ and integrate the SDE (7) using the Euler–Maruyama method. To approximate the Koopman operator for this system, we use $n=25$ Gaussian functions in the domain $[0,2]\times[-2,2]$ as shown in Figure 4b. That is, the centers of all the basis functions are initially in the right half of the domain. The optimized basis functions are shown in Figure 4c. The centers moved and also the bandwidths changed significantly, leading to a more accurate approximation of the Koopman operator. The second eigenfunction of the approximated Koopman operator, separating the two deeper wells, is shown in Figure 4d.
@@ 4.2.1 Chua’s circuit
Chua’s circuit [34, 35] is given by
$\begin{split}\dot{x}_{1}&=\alpha[x_{2}-x_{1}-f(x)],\\ \dot{x}_{2}&=\frac{1}{RC_{2}}[x_{1}-x_{2}+Rz],\\ \dot{x}_{3}&=-\beta x_{2},\end{split}$
where $f(x)=-b\sin\big{(}\frac{\pi x_{1}(t)}{a}+d\big{)}$ describes the electrical response of the resistor, see [36]. We generate training data using $a=2.6,b=0.11,c=7,d=0,\alpha=10.2,\beta=14.286$, and $R=C_{2}=1$. To illustrate how the parameters affect the dynamics, we plot the reconstruction error for varying values of $w_{1}$ using the basis functions
@@
The results are shown in Figure 5a. The minimum reconstruction error is obtained for $w_{1}\approx 1.208$, i.e., $a\approx 2.6$, which is the actual value of the parameter. The governing equations cannot be identified using standard SINDy, unless we explicitly choose the basis function $\sin(\frac{\pi}{2.6}\hskip 1.00006ptx_{1})$. We use the proposed framework with the above set of parametric basis functions $\mathcal{D}$ to find the optimal matrix $\Xi$ and parameters $w$. The convergence of the algorithms is illustrated in Figures 5b and 5c, respectively. The dynamics of the discovered equations in Figure 5d show that the algorithm determines the correct matrix $\Xi$ and the correct value of the parameter $w_{1}$.
Figure 5: Results for Chua’s circuit. (a) SINDy reconstruction error as a function of $w_{1}$. (b) & (c) Convergence of the approximation errors for the loss function for $\Xi$ and $w$, respectively. (d) Reconstruction of the dynamics using the optimization algorithms. The orange graph completely overlaps the blue, representing a perfect reconstruction of the dynamics. The green trajectory shows the resulting dynamics using a different value of $w_{1}$.
@@ 4.2.2 Nonlinear heat PDE
$\Theta(U(\chi))=\big{[}1,U,U_{x},UU_{x},U^{2}U_{x},UU_{xx},U^{2}U_{xx},e^{\chi U}U_{x}^{2},e^{\chi U}U_{xx}\big{]}.$
We apply the alternating optimization to the reconstruction error (6) with $\epsilon=10^{-4}$. The algorithm gives us the solution vector $\xi$ containing the coefficients for the different basis functions and the parameter $\chi=-0.76$. Using iterative hard thresholding, we obtain
$u_{t}=-0.075\hskip 1.00006pte^{\chi u}u_{x}^{2}+0.055\hskip 1.00006pte^{\chi u}u_{xx}.$
The numerical errors in the coefficients of the identified terms are caused by the finite difference approximations of the derivatives. Additionally, the gradient descent algorithm might not have converged to the global minimum. However, the example shows that our method allows us to identify parameter-dependent PDEs from data.
@@ 5 Conclusion and future work
There are, however, open problems. One of the major challenges is selecting proper initial conditions and learning rates for the gradient descent algorithms. The algorithms might diverge or get stuck in a local minimum if the initial conditions or step sizes are poorly chosen. This leads to several interesting directions to extend this work in the future: A theoretical analysis of the optimization problems would help us understand the loss functions and the existence of local minima, which might allow us to choose suitable initial conditions and step sizes. An extension of this work would be to consider kernel-based variants of the data-driven algorithms, such as kernel EDMD (kEDMD) [9, 10]. The proposed framework could then be used to optimize the kernel parameters.
@@ Acknowledgments
MT was supported by the EPSRC Centre for Doctoral Training in Mathematical Modeling, Analysis and Computation (MAC-MIGS) funded by the UK Engineering and Physical Sciences Research Council (grant EP/S023291/1), Heriot-Watt University and the University of Edinburgh. NKC is supported by an EPSRC-UKRI AI for Net Zero Grant: “Enabling CO2 Capture and Storage Projects Using AI”, (Grant EP/Y006143/1). NKC is also supported by a CityU startup grant.

### v2

@@ 4.1.3 Triple-well 2D
The potential is visualized in Figure 4a. We set $\beta=1.68$ and uniformly sample $m=100\hskip 1.00006pt000$ training data points in the domain $[-2,2]\times[-2,2]$ and integrate the SDE (7) using the Euler–Maruyama method. To approximate the Koopman operator, we use $n=25$ Gaussian functions in the domain $[0,2]\times[-2,2]$ as shown in Figure 4b. That is, the centers of all the basis functions are initially in the right half of the domain. The optimized basis functions are shown in Figure 4c. The centers moved and also the bandwidths changed significantly, leading to a more accurate approximation of the Koopman operator. The second eigenfunction of the approximated Koopman operator, separating the two deeper wells, is shown in Figure 4d.
@@ 4.2.1 Chua’s circuit
We consider the modified Chua circuit [37] given by
$\begin{split}\dot{x}_{1}&=\alpha[x_{2}-f(x)],\\ \dot{x}_{2}&=x_{1}-x_{2}+x_{3},\\ \dot{x}_{3}&=-\beta x_{2},\end{split}$
where $f(x)=-b\sin\big{(}\frac{\pi x_{1}(t)}{a}+d\big{)}$ describes the electrical response of the resistor, see [38, 39]. We generate training data using $a=2.6,b=0.11,d=0,\alpha=10.2,\beta=14.286$. To illustrate how the parameters affect the dynamics, we plot the reconstruction error for varying values of $w_{1}$ using the basis functions
@@
The results are shown in Figure 5a. The minimum reconstruction error is obtained for $w_{1}\approx 1.208$, i.e., $a\approx 2.6$, which is the correct value of the parameter. We see that there are two regions in the energy landscape. Selecting the initial value for $w_{1}$ in Region 1, we obtain the global minimum, i.e., zero reconstruction error, whereas choosing the initial value to be in Region 2, the algorithm converges to the local minimum. The governing equations cannot be identified using standard SINDy, unless we explicitly choose the basis function $\sin(\frac{\pi}{2.6}\hskip 1.00006ptx_{1})$. We use the proposed framework with the above set of parametric basis functions $\mathcal{D}$ to find the optimal matrix $\Xi$ and parameters $w$. The convergence of the algorithms is illustrated in Figures 5b and 5c, respectively. The dynamics of the discovered equations shown in Figure 5d illustrate that the algorithm determines the correct matrix $\Xi$ and the correct value of the parameter $w_{1}$, provided that we start in Region 1.
Figure 5: Results for Chua’s circuit. (a) SINDy reconstruction error as a function of $w_{1}$. (b) & (c) Convergence of the approximation errors for the loss functions for $\Xi$ and $w$, respectively. (d) Reconstruction of the dynamics using the optimization algorithms. The orange graph completely overlaps the blue, representing a perfect reconstruction of the dynamics. The green trajectory shows the resulting dynamics using a different value of $w_{1}$.
@@ 4.2.2 Nonlinear heat PDE
$\Theta(U(\chi))=\big{[}1,U,U_{x},UU_{x},U^{2}U_{x},UU_{xx},U^{2}U_{xx},e^{\chi U}U_{x}^{2},e^{\chi U}U_{xx}\big{]}$
and apply the alternating optimization algorithm to the reconstruction error (6). The resulting vector $\xi$ then contains the coefficients for the different terms. We obtain $\chi=-1.095$ and the PDE
$u_{t}=-0.114\hskip 1.00006pte^{\chi u}u_{x}^{2}+0.105\hskip 1.00006pte^{\chi u}u_{xx}.$
The small difference between the true PDE and the identified PDE is caused by the finite difference approximations of the derivatives. Nevertheless, the example shows that our method allows us to identify parameter-dependent PDEs from data.
Figure 6: Simulation of the nonlinear heat equation. (a) Surface plot of the solution of the PDE. (b) Contour plot of the error between the original and the discovered PDE. (c) Convergence of the algorithm to the correct value of $\chi$.
@@ 5 Conclusion and future work
There are, however, open problems. One of the major challenges is selecting proper initial conditions and learning rates for the gradient descent algorithms. The algorithms might diverge or get stuck in local minima if the initial conditions or step sizes are chosen poorly. This leads to several interesting directions to extend this work in the future: A theoretical analysis of the optimization problems would help us understand the loss functions and the existence of local minima, which might allow us to choose suitable initial conditions and step sizes. An extension of this work would be to consider kernel-based variants of the data-driven algorithms, such as kernel EDMD [9, 10]. The proposed framework could then be used to optimize the kernel parameters.
@@ Acknowledgments
MT was supported by the EPSRC Centre for Doctoral Training in Mathematical Modeling, Analysis and Computation (MAC-MIGS) funded by the UK Engineering and Physical Sciences Research Council (grant EP/S023291/1), Heriot–Watt University and the University of Edinburgh. NKC is supported by an EPSRC-UKRI AI for Net Zero Grant: “Enabling CO2 Capture and Storage Projects Using AI”, (Grant EP/Y006143/1). NKC is also supported by a City University of Hong Kong Start-up Grant, project number 7200809. We would like to thank D.E. Shaw Research for providing the Chignolin data.